OpenStudy (anonymous):

@jim_thompson5910 can you help me please write the explicit general term for the arithmetic sequence where a2 =4 and a6=32 and find the 25 term.

jimthompson5910 (jim_thompson5910):

an = a1*(r)^(n-1) a2 = a1*(r)^(2-1) a2 = a1*r^1 a2 = a1*r

jimthompson5910 (jim_thompson5910):

a2 = a1*r 4 = a1*r a1 = 4/r

jimthompson5910 (jim_thompson5910):

an = a1*(r)^(n-1) a6 = (4/r)*(r)^(n-1) a6 = 4*(r)^(6-2) 32 = 4*(r)^(4) 32/4 = r^4 8 = r^4 r^4 = 8 r = (8)^(1/4) r = 1.68179 a1 = 4/r a1 = 4/1.68179 a1 = 2.3784182

jimthompson5910 (jim_thompson5910):

So after all this, we can say that the sequence is roughly an = 2.3784182(1.68179)^(n-1)

jimthompson5910 (jim_thompson5910):

Your last task is to plug in n = 6 and simplify to get a6

jimthompson5910 (jim_thompson5910):

oh wait, i'm thinking of something else you plug in n = 25 and simplify to find a25

OpenStudy (anonymous):

okaii and that will be my final answer

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thank you once again :)

jimthompson5910 (jim_thompson5910):

yw